Chapter 13, Problem 045 A certain satellite travels in an approximately circular orbit of radius 7.7 × 106 m with a period of 6 h 36 min. Calculate the mass of its planet from this information. Number the tolerance is +/-296 Click if you would like to Show Work for this question: pShow.Work Units LINK TO
Answer #1

Given that Radius of the circular orbit r = 7.7 × 106 m Time period T = 6 h 36 min = 6x3600 s + 36 × 60 s = 23760 s Now from